OpenStudy (paki):

Find the value of "C"...? 5C+8 and 12C-3

OpenStudy (kc_kennylau):

Is this the whole question?

OpenStudy (kc_kennylau):

Is the word "and" considered to be equivalent to the word "equal"?

OpenStudy (nincompoop):

sighs...

OpenStudy (anonymous):

if its 2 separate questions try 5c = - 8 c = -8/5 and 12c = 3 c = 3/12